The state highway department is studying traffic patterns on one of the busiest highways in the state. As part of the study, the department needs to estimate the average number of vehicles that pass an intersection each day. A random sample of 64 days gives us a sample mean of 14,205 cars and a sample standard deviation of 1,010 cars. What is the 92% confidence interval estimate of
Answer #1


Answer Date: 29/12/2019 The 92% confidence interval estimate of the average number of vehicles that pass an intersection each92%C.I. = # It, xs/In = 14205 +1.779x1010/164 = 14205 + 224.65 = 13980.35 to 14429.65 The 92% confidence interval estimate of